OpenStudy (anonymous):

rewrite 3rd root square root 27x-81-5

OpenStudy (anonymous):

you mean \[\sqrt[3]{\sqrt{27x-81-5}}\]? i doubt that is the question

OpenStudy (anonymous):

seems more likely that it is \[\sqrt[3]{27x-81}\]

OpenStudy (anonymous):

but i am not sure, so let me know

OpenStudy (anonymous):

yes your right